OpenStudy (anonymous):

Tell what kind of graph this equation yields, and identify the domain and range. 4x^2+4y^2=16

OpenStudy (ijlal):

are you familiar with circle equations?

OpenStudy (ijlal):

Equation of a circle is \[(x-a)^2+(y-b)^2=r^2\] where (a,b) is the center of the circle and r is the radius so lets see the equation in the question shall we it is \[4x^2+4y^2=16\] we can also write this as \[4(x-0)^2+4(y-0)^2=4^2\] if we compare (a.b) it shows us that (a,b)-->(0,0) that is the origin so the center of the circle is an origin and radius is 4 if we simplify the equation lets say divide the equation by 4 we get \[x^2+y^2=2^2\] this means the radius is 2

OpenStudy (ijlal):

let us sketch this

OpenStudy (ijlal):